a motorboat travels 60 kms upstream in 10 hrs if it returns in 8 hrs. . Find th rate of the current and the boat rates are constant during the trip.
Answer by venugopalramana(3286)

a motorboat travels 60 kms upstream in 10 hrs if it returns in 8 hrs. . Find th rate of the current and the boat rates are constant during the trip.
LET SPEED OF BOAT IN STILL WATER=B
SPEED OF WATER=W
UP STREAM SPEED =B-W
T=10 HRS....D=60 KM
10(B-W)=60
B-W=6................I
DOWN STREAM SPEED = B+W
T=8 HRS.....D=60 KM
8(B+W)=60
B+W=60/8=7.5.................II
EQN.I+EQN.II
2B=13.5
B=6.75 KMPH
W=7.5-6.75=0.75 KMPH